### v2

Abstract
Accurate and efficient computation of Floquet multipliers and subspaces is essential for analyzing limit cycle in dynamical systems and periodic steady state in Radio Frequency simulation. This problem is typically addressed by solving a periodic linear eigenvalue problem, which is discretized from the linear time-periodic system using one-step collocation methods. Collocation methods become costly for large-scale cases. Our alternative approach is to use multistep methods. The multistep method leads to a periodic polynomial eigenvalue problem (pPEP), and introduces additional parasitic periodic eigenvalues. We prove that as the stepsize decreases, the computed Floquet multipliers and their associated invariant subspace converge with higher order, while the parasitic periodic eigenvalues converge to zero geometrically and therefore Floquet multipliers are not affected by those parasitic ones. A memory-efficient algorithm pTOAR is designed to solve the large-scale pPEP. Its computational and memory costs are almost independent of the choice of multistep methods. Numerical results coincide with our convergence analysis, and also demonstrate the efficiency of pTOAR.
keywords:
Floquet Multiplier, Multistep Method, Periodic Polynomial Eigenvalue Problem, Subspace Perturbation
{MSCcodes}
65L15, 65F15, 65L70
1 Introduction
A linear time-periodic (LTP) system is described by
(1) $\frac{\mathrm{d}x(t)}{\mathrm{d}t}=G(t)x(t),$
where $G(t)\in\mathbb{R}^{n\times n}$ is a $T$-periodic matrix function, i.e., $G(t+T)=G(t)$, and $x(t)\in\mathbb{R}^{n}$ is the state vector. The monodromy matrix of Eq.˜1 is obtained by integrating Eq.˜1 from 0 to $T$ with initial value matrix $X(0)=I_{n}$. The Floquet multipliers and eigenvectors are the eigenpairs of the monodromy matrix. Efficient and accurate computation of Floquet multipliers and corresponding eigenvectors is essential for analyzing LTP systems.
The LTP systems find applications in a wide range of areas, including dynamic simulation [1], communications [15], control theory [6, 4] and Radio Frequency circuit simulation [9]. The central problem in these fields is the stability analysis for the limit cycles of nonlinear autonomous systems. The stability is completely characterized by the Floquet multipliers and eigenvectors of the variational equation along the limit cycle in the form of Eq.˜1 [20].
The computation of Floquet multipliers and eigenvectors generally involves two key steps: discretization of Eq.˜1 and solving the resulting eigenvalue problem. The simplest implementation is to use backward Euler (BE) method to obtain a periodic eigenvalue problem (pEVP), which can be solved by periodic QR algorithm [5, 18]. Existing high-precision software packages such as AUTO-07p [12] and MATCONT [11] adopt high-order one-step collocation methods for discretization. Collocation methods require extra function evaluations at Gaussian collocation points within each subinterval, and a condensation step to eliminate the extra collocation points. This framework makes collocation well-suited for modest-scale cases whose coefficient function $G(t)$ is given by an analytical expression. For large-scale systems, collocation methods become computationally expensive and memory-intensive [20], particularly as the condensation may break the inherent sparsity of the systems. In some practical cases, collocation methods also struggle if extra function evaluations are not permitted.
This paper applies multistep methods to discretize Eq.˜1 that induces a periodic polynomial eigenvalue problem (pPEP). Multistep methods improve accuracy under the same sampling used by BE. However, applying a $d$-step method, the resulting pPEP possesses $nd$ periodic eigenpairs under cyclic symmetry; and the pPEP is much more complicated to solve.
By introducing parasitic LTP systems, we prove that n(d1)n(d-1) periodic eigenvalues decay to 0 at rate $\mathcal{O}(\Delta t)$ as the maximum stepsize $\Delta t$ goes to 0. The remaining $n$ eigenpairs converge to Floquet eigenpairs at rate $\mathcal{O}(\Delta t^{s})$, where $s$ is the consistency order of the multistep method. Therefore, we only need to solve the $n$ dominant periodic eigenvalues of pPEP.
To efficiently solve the eigenpairs of the pPEP, we propose the pTOAR. In particular, we show that using high-order multistep method incurs only negligible additional cost compared to the first-order one-step backward Euler method.
The remainder of this paper is organized as follows. In Section˜2, the necessary background on Floquet theory, one-step collocation methods, multistep methods, and subspace perturbation theory are reviewed. In Section˜3, we discuss the formulation of pPEP. In Section˜4, the convergence results are established for the eigenpairs using tools from subspace perturbation theory. In Section˜5, we present the pTOAR for solving pPEP, together with an analysis of its computational and memory costs. Numerical experiments are reported in Section˜6, and conclusions are drawn in Section˜7.
2 Preliminaries
2.1 Floquet Theory
Floquet theory provides the structure of the solutions of LTP systems Eq.˜1. We recall the key results of Floquet theory, and the comprehensive version can be found in [2, Chapter 5] and [14, Chapter 2].
For LTP system Eq.˜1 with period $T$, let $X(t)$ denote a fundamental solution matrix, defined by $\dot{X}(t)=G(t)X(t)$, where the initial value matrix $X(0)=X_{0}$ is any nonsingular matrix. Define the state transition matrix $\Phi(t,s)$ by as the mapping from the principal fundamental solution matrix at time $s$ to that at time $t$, that $\Phi(t,s)X(s)=X(t)$.
Theorem 2.1 (Floquet Representation Theorem [2][Theorem 20.9]).
For LTP system Eq.˜1, there exists a nonsingular constant matrix $M$ and a $T$-periodic matrix function $U(t)\in C^{1}(\mathbb{R},\mathbb{C}^{n^{2}})$ such that its state transition matrix $\Phi(t,0)$ can be written as
(2) $\Phi(t,0)=U(t)e^{Mt}.$
Thus, the state transition matrix over a period $\Phi(t+T,t)$ can be decomposed by:
(3) $\Phi(t+T,t)=U(t)e^{MT}U^{-1}(t),$
Since $M$ is a constant matrix, and $T$ is the period, all matrices $\Phi(t+T,t)$ are similar. The eigenvalues of $M$ and $e^{MT}$ are called Floquet exponent and Floquet multipliers. For simplicity, we assume all Floquet multipliers are semi-simple. Thus, we assume $e^{MT}$ is diagonal, and Eq.˜3 becomes an eigenvalue decomposition of $\Phi(t+T,t)$. The columns of $U(t)$ and the rows of $U^{-\mathrm{H}}(t)$ provide the corresponding right and left Floquet eigenfunctions, respectively.
To obtain the Floquet multipliers, we only need to consider the eigenvalue problem Eq.˜4, which corresponds to Eq.˜3 at $t=0$:
(4) $\Phi(T,0)u(0)=u(0)e^{\mu T}.$
$u(0)$ is called the Floquet eigenvector associated with the Floquet multiplier $\lambda=e^{\mu T}$.
Though Eq.˜4 is only a slice of Eq.˜3 at $t=0$, it is sufficient to solve Floquet eigenfunctions $U(t)$ within the period. Once all the Floquet multipliers $e^{MT}$ are obtained, we can figure out all the Floquet exponents $M$, and thus solve Floquet eigenfunctions $U(t)$ from Eq.˜2. We should also note that $u(0)$ in Eq.˜4 is the value of the Floquet eigenfunction at $t=0$.
2.2 One-step Methods for Solving Floquet Eigenpairs
Let the period $[0,T]$ be partitioned at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$, with stepsizes $h_{i}=t_{i}-t_{i-1}$ allowed to be variable, and with maximum $h$. Rewrite Eq.˜3 on the time grid:
$\Phi(t_{p},t_{p-1})\cdots\Phi(t_{2},t_{1})\Phi(t_{1},t_{0})u(t_{0})=u(t_{0})e^{\mu T}.$
When $h$ is small enough, one-step methods approximate each $\Phi(t_{i},t_{i-1})$ by a matrix $A^{(i)}$ for $i=1,\ldots,p$, and approximate the eigenpair $\left(u(t_{0}),e^{\mu T}\right)$ by $\left(u_{0},\lambda\right)$. This leads to a periodic eigenvalue problem [18]:
(5) $A^{(p)}\cdots A^{(2)}A^{(1)}u_{0}=u_{0}\lambda.$
One of the simplest one-step method is backward Euler (BE). BE only requires function evaluation of $G(t_{i})$ at time $t_{i}$, and $A^{(i)}=\left[I-G(t_{i})h_{i}\right]^{-1}$ when applied to Eq.˜1. However, its first-order accuracy is usually insufficient in practical.
High-order one-step methods, such as collocation, improve discretization accuracy. In what follows, we briefly outline this procedure; a comprehensive process can be found in [20].
The first step is to choose the number of collocation points $k$ within an interval $t_{i-1},t_{i}]$, where a larger $k$ corresponds higher local approximation accuracy. Collocation methods introduce internal stage variables at $k$ equidistant nodes and constructs a Lagrange polynomial interpolant through these points to approximate the solution $x(t)$. The interpolant is then required to satisfy Eq.˜1 at $k$ Gaussian collocation points within $t_{i-1},t_{i}]$, which needs extra function evaluation of the system matrix $G(t)$ at these additional Gaussian collocation points.
The next step, known as condensation, is eliminating the intermediate stage variables to obtain the one-step relation $x_{i}=A^{(i)}x_{i-1}$, typically via an LU elimination [20]. This process yields an explicit construction of the local state transition matrix $A^{(i)}$.
Collocation methods achieve higher accuracy than BE, and are efficient for cases that $G(t)$ is analytically known. It is also adopted by existing mature public software packages for computing and analyzing limit cycles, such as MATCONT [11] and AUTO-07p [12]. However, if $G(t)$ is only available on given time grids, the collocation methods become both impractical and costly:
  1. 1.
    they rely on temporal internal function evaluation on Gaussian points of each interval for accuracy, which is expensive and often infeasible in large-scale applications;
  2. 2.
    the condensation procedure incurs significant computational cost, often about twice that of a direct eigendecomposition [20];
  3. 3.
    the condensed transition matrix $A^{(i)}$ is typically dense and must be formed explicitly, increasing memory usage and limiting the advantage of matrix-free solvers.
2.3 Multistep Methods for Solving Initial Value Problems
Multistep methods offer an alternative approach to enhance discretization accuracy by utilizing previously computed values at neighboring grid points. We review their general form and recall standard notions of consistency and stability [17]. These results are essential for solving initial value problems, and also provide the foundation for our later analysis of Floquet multipliers using multistep discretizations.
Definition 2.2 (Variable Stepsize Multistep Methods [17, Chapter 3.5]).
Let $0=t_{0}<\cdots<t_{p}=T$ be a time grid on the interval $[0,T]$. We define the stepsize $h_{i}=t_{i}-t_{i-1}$ and the stepsize ratio $\omega_{i}=h_{i}/h_{i-1}$. We define the ratio vector $\boldsymbol{\omega}_{i}\in\mathbb{R}^{d}$ as: $\boldsymbol{\omega}_{i}=(\omega_{i-d+1},\ldots,\omega_{i}).$ For an ODE $\dot{x}=f(t,x)$, a $d$-step method updates the solution at $t_{i}$ according to:
(6) $\sum_{j=0}^{d}\alpha_{j}^{(i)}x_{i-d+j}=h_{i}\sum_{j=0}^{d}\beta_{j}^{(i)}f(t_{i-d+j},x_{i-d+j}),$
where the coefficients are determined by smooth functions of the local step ratios, i.e., $\alpha_{j,i}=\alpha_{j}(\boldsymbol{\omega}_{i})$ and $\beta_{j,i}=\beta_{j}(\boldsymbol{\omega}_{i})$, and $\left\lvert\alpha_{0,i}\right\rvert+\left\lvert\beta_{0,i}\right\rvert\neq 0$ for irreducibility condition.
For detailed derivations of these coefficients, see [17, Chapter 3]. As an illustration, the second-order BDF method (Gear2) has coefficients given by Eq.˜7.
(7) $\alpha_{0,i}=\frac{{\omega_{i}}^{2}}{1+2\omega_{i}},\,\alpha_{1,i}=-1-\alpha_{0,i},\alpha_{2,i}=1,\,\beta_{2,i}=\frac{1+\omega_{i}}{1+2\omega_{i}},\,\beta_{1,i}=\beta_{0,i}=0.$
We now turn to the fundamental notions of consistency and stability for such multistep methods.
Definition 2.3 (Consistency [17][Chapter 3, 5.17]).
A multistep method Eq.˜6 is consistent of order $s$ if, when applied to a sufficiently smooth function $x(t)$, its local truncation error $E_{i}$ satisfies $\left\lVert E_{i}\right\rVert=\mathcal{O}(h_{i}^{s+1})$ as $h_{i}\to 0$, where
(8) $E_{i}=\sum_{j=0}^{d}\alpha_{j,i}x(t_{i-d+j})-h_{i}\sum_{j=0}^{d}\beta_{j,i}\frac{\mathrm{d}x}{\mathrm{d}t}(t_{i-d+j}).$
The stability of a multistep method is to guarantee boundedness of the numerical solution. The trivial ODE $\dot{x}=0$ is usually chosen to be the test equation. Apply the multistep discretization, we obtain the homogeneous recurrence $\sum_{j=0}^{d}\alpha_{j,i}x_{i-d+j}=0.$ Now, we introduce $X_{i}=(x_{i-d+1},\cdots,x_{i-1},x_{i})^{\mathrm{T}}$ and the companion matrix
(9) $L_{\alpha;i}=\begin{bmatrix}&1&&\\ &&\ddots&\\ &&&1\\ -\alpha_{0,i}&-\alpha_{1,i}&\cdots&-\alpha_{d-1,i}\end{bmatrix},\,F_{\alpha;i+j,i}\equiv L_{\alpha;i+j}\cdots L_{\alpha;i+2}L_{\alpha;i+1}.$
Definition 2.4 (Stability [17][Chapter 3, Definition 5.4]).
The multistep method in Eq.˜6 is called stable if for all $i$ and $j\geq 0$, there exists a constant $C$, such that
(10) $\left\lVert F_{\alpha;i+j,i}\right\rVert\leq C.$
For a fixed multistep method, the matrices $F_{\alpha;i+j,i}$ in Definition˜2.4 depend on the chosen time grid, because the coefficients $\alpha_{j,i}$ and $\beta_{j,i}$ are functions of the local stepsize ratios. As a result, stability can only be verified for each specific grid individually, which becomes cumbersome when the stepsizes vary. [17] shows that, if the method is stable under constant stepsizes and certain additional conditions are satisfied, there exist constants $0<\omega<1<\Omega$ such that, whenever the step ratios satisfy $\omega_{i}\in[\omega,\Omega]$, the multistep method remains stable on all corresponding grids. For simplicity, throughout this paper we assume that the method is stable on the chosen variable stepsize grid.
3 pPEP Derived by Multistep Discretization of LTP System
When a multistep method is applied to an LTP system Eq.˜1, it gives rise to a linear difference equation with periodic coefficient matrices, referred to as a periodic difference equation. In this section, we will show that this periodic difference equation corresponds to a periodic polynomial eigenvalue problem (pPEP). Furthermore, the pPEP admits equivalent representations either as a polynomial eigenvalue problem with block-cyclic coefficient matrices or as a periodic eigenvalue problem with coefficients taking the companion form of the underlying difference equation.
3.1 Derivation of pPEP and its Linearization
Consider a partition of one period $[0,T]$ at time points $0=t_{0}<t_{1}<\cdots<t_{p}=T$. This grid is extended to an infinite sequence $\left\{t_{i}\right\}_{i\in\mathbb{Z}}$ satisfying $t_{i+p}=t_{i}+T$, which defines a periodic time grid. Throughout this paper, quantities defined on this periodic time grid that exhibit $p$-periodicity are indexed with a superscript $(i)$ to indicate their cyclic dependence. Accordingly, the multistep coefficients $\left(\alpha_{j,i},\beta_{j,i}\right)$ are denoted by $\left(\alpha_{j}^{(i)},\beta_{j}^{(i)}\right)$, the stepsizes by $h^{(i)}=t_{i}-t_{i-1}$, the maximum step by $h=\max_{i=1}^{p}h^{(i)}$, and the step ratio by $\omega^{(i)}$.
Applying a linear $d$-step method to Eq.˜1 yields a linear $d$-order difference equation of the form
(11) $\sum_{j=0}^{d}A_{j}^{(i)}x_{i-d+j}=0,$
where the coefficient matrices $A_{j}^{(i)}\in\mathbb{C}^{n\times n}$ are $p$-periodic in $i$ and given by
(12) $A_{j}^{(i)}=\alpha_{j}^{(i)}I_{n}-\beta_{j}^{(i)}G^{(i-d+j)}h^{(i)},\quad j=0,\ldots,d.$
To compute Floquet multipliers, we seek an eigenvalue formulation associated with Eq.˜11. Motivated by the discrete Floquet theorem for one-step periodic recursions [7], we substitute the Floquet-type Ansatz $x_{i}=u^{(i)}\theta^{i}$ into the general $d$-step recursion Eq.˜11, where $u^{(i)}$ is $p$-periodic. This leads to a periodic polynomial eigenvalue problem characterizing the multistep dynamics.
Definition 3.1 (Periodic Polynomial Eigenvalue Problem (pPEP)).
Let $\{A_{j}^{(i)}\}_{j=0,\ldots,d}^{i=1,\ldots,p}$ denote $d+1$ sequences of matrices in $\mathbb{C}^{n\times n}$, each of period $p$. The periodic polynomial eigenvalue problem (pPEP) consists of finding a scalar $\theta\in\mathbb{C}$ and a nonzero $p$-periodic sequence $\{u^{(i)}\}_{i\in\mathbb{Z}}$, $u^{(i)}\subset\mathbb{C}^{n}$ such that
(13) $\sum_{j=0}^{d}A_{j}^{(i)}\,u^{(i-d+j)}\theta^{j}=0,\qquad i=1,\ldots,p,$
where the indices $(i)$ in the superscript are understood $i$ modulo $p$.
Stacking $p$ components in Eq.˜13 together leads to a polynomial eigenvalue problem (PEP) Eq.˜14 of degree $d$ with block-cyclic coefficient matrices to reflect the periodicity. To illustrate this, we define the block-diagonal matrices $\mathcal{A}_{j}=\operatorname{diag}_{i}(A_{j}^{(i)}),\,j=0,\ldots,d,$ and the cyclic shift operator $\mathcal{C}=C_{p}\otimes I_{n}$, $C_{p}=[e_{2},\dots,e_{p},e_{1}],$ where $e_{i}$ denotes the $i$-th standard basis vector of $\mathbb{R}^{p}$.
(14) $\left(\sum_{j=0}^{d}\mathcal{A}_{j}\mathcal{C}^{d-j}\theta^{j}\right)\operatorname{col}(u^{(i)})_{i=1}^{p}=0,\quad\operatorname{col}(u^{(i)})_{i=1}^{p}\equiv\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ \vdots\\ u^{(p)}\end{bmatrix}\in\mathbb{C}^{np},$
A simple example monic pPEP for $d=2$ and $p=3$ is given by Eq.˜15:
(15) $\left(I_{3n}\theta^{2}+\begin{bmatrix}&&A_{1}^{(1)}\\ A_{1}^{(2)}&&\\ &A_{1}^{(3)}&\\ \end{bmatrix}\theta+\begin{bmatrix}&A_{0}^{(1)}&\\ &&A_{0}^{(2)}\\ A_{0}^{(3)}&&\\ \end{bmatrix}\right)\begin{bmatrix}u^{(1)}\\ u^{(2)}\\ u^{(3)}\\ \end{bmatrix}=0.$
If $(\theta,\operatorname{col}\{u^{(i)}\}_{i=1}^{p})$ is an eigenpair, then for each $k=0,\ldots,p-1$,
$\left(\theta\zeta_{p}^{k},\operatorname{col}\{u^{(i)}\zeta_{p}^{ik-k}\}_{i=1}^{p}\right)$
is also an eigenpair, where $\zeta_{p}=e^{\jmath 2\pi/p}$ is a primitive $p$-th root of unity, that $\jmath^{2}=-1$. Among these, the periodic eigenvalue is defined as $\lambda=\theta^{p}$, and the periodic eigenvector as the sequence $\operatorname{col}\{u^{(i)}\}_{i=1}^{p}$ corresponding to the root $\theta$ with minimal argument $\arg(\theta)\in[0,2\pi/p)$, the periodic eigenpair is referred to as $(\theta^{p},\operatorname{col}\{u^{(i)}\}_{i=1}^{p})$.
The pPEP is called monic if its leading coefficient matrices satisfy $A_{d}^{(i)}=I_{n}$ for all $i=1,\ldots,p$. For pPEPs arising from multistep discretizations of LTP ODE systems, this assumption entails no loss of generality. For sufficiently small step sizes $h^{(i)}$, the matrices $A_{d}^{(i)}$ defined in Eq.˜12 are nonsingular. Thus, by left-multiplying the $i$-th equation in Eq.˜13 by $\left(A_{d}^{(i)}\right)^{-1}$, the pPEP can always be transformed into an equivalent monic form.
The linearization of the PEP Eq.˜14 is
(16) $\begin{bmatrix}&I_{np}&&\\ &&\ddots&\\ &&&I_{np}\\ -\mathcal{A}_{0}\mathcal{C}^{d}&-\mathcal{A}_{1}\mathcal{C}^{d-1}&\cdots&-\mathcal{A}_{d-1}\mathcal{C}\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}\theta.$
3.2 Linearized pPEP as a Periodic Eigenvalue Problem
The pPEP Eq.˜14 in PEP form can be transformed into a periodic eigenvalue problem (pEVP) of dimension $nd$ and period $p$. The transformation is obtained by a linearization of PEP followed by a permutation.
The eigenvector in Eq.˜16 exhibits a two-level structure: the inner structure encodes the periodicity of the coefficients, while the outer structure arises from the linearization of Eq.˜14. To make this structure explicit, we introduce a permutation matrix $\mathcal{P}$ that reorders the stacked vector:
$\mathcal{P}\begin{bmatrix}\operatorname{col}(u^{(i)})_{i=1}^{p}\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta\\ \vdots\\ \operatorname{col}(u^{(i)})_{i=1}^{p}\theta^{d-1}\end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix},\,\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\begin{bmatrix}u^{(i-d+1)}\\ u^{(i-d+2)}\theta\\ \vdots\\ u^{(i)}\theta^{d-1}\\ \end{bmatrix}$.
After this permutation, Eq.˜16 takes the following block-cyclic form:
(17) $\begin{bmatrix}&&&L^{(1)}\\ L^{(2)}&&&\\ &\ddots&&\\ &&L^{(p)}&\\ \end{bmatrix}\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}=\begin{bmatrix}\operatorname{col}(u^{(1-j)}\theta^{d-j})_{j=1}^{d}\\ \operatorname{col}(u^{(2-j)}\theta^{d-j})_{j=1}^{d}\\ \vdots\\ \operatorname{col}(u^{(p-j)}\theta^{d-j})_{j=1}^{d}\\ \end{bmatrix}\theta,$
where each block $L^{(i)}$ is the companion matrix of Eq.˜11 at step $i$:
(18) $L^{(i)}=\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ -A_{0}^{(i)}&-A_{1}^{(i)}&\cdots&-A_{d-1}^{(i)}\\ \end{bmatrix}$.
The block-cyclic structure implies that, for every $i=1,\ldots,p$, Eq.˜17 reduces to $p$ periodic eigenvalue problems of dimension $nd$:
(19) $L^{(i+p)}L^{(i+p-1)}\cdots L^{(i+1)}\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}=\operatorname{col}(u^{(i-j)}\theta^{d-j})_{j=1}^{d}\theta^{p},\,i=1,\ldots,p.$
4 Convergence Analysis of Solving Floquet eigenpairs from pPEP
In this section, we analyze whether and in what sense the periodic eigenpairs obtained from the linearized form of the pPEP Eq.˜19 approximate the Floquet eigenpairs of the underlying LTP system, and characterize the role of the additional eigenvalues introduced by the multistep discretization.
4.1 Connection between Periodic Eigenpairs of pPEP and Floquet Eigenpairs
This subsection motivates the subsequent convergence analysis by examining the structural relationship between the linearized pPEP Eq.˜14 and the continuous Floquet eigenvalue problem Eq.˜4.
Each companion matrix $L^{(i)}$ in Eq.˜18 performs one-step propagation of the $d$-stacked state vectors of Eq.˜11 [16]:
(20) $\operatorname{col}(x_{j})_{j=i-d+1}^{i}=L^{(i)}\operatorname{col}(x_{j})_{j=i-d}^{i-1}.$
Consequently, the discrete stacked propagator of Eq.˜11 from step $i$ to $i+j$ is
(21) $F_{i+j,i}=L^{(i+j)}L^{(i+j-1)}\cdots L^{(i+1)},$
and in particular $F_{i+p,i}$ represents the one-period discrete stacked propagator. Because $L^{(i)}$ is $p$-periodic, $F_{i+p,i}$ for all $i\in\mathbb{Z}$ satisfies similarity relation $F_{i+p,i}=F_{i,0}F_{p,0}F_{i,0}^{-1},$ so it suffices to consider $i=0$. The propagation from step 0 to $p$ is given by
(22) $\operatorname{col}(x_{j})_{j=p-d+1}^{p}=F_{p,0}\,\operatorname{col}(x_{j})_{j=-d+1}^{0}.$
and the pPEP Eq.˜19 amounts to finding a sequence $x_{i}$ satisfying $x_{i+p}=x_{i}\lambda,$ i.e. $\operatorname{col}(x_{j})_{j=p-d+1}^{p}=\operatorname{col}(x_{j})_{j=-d+1}^{0}\lambda$
By comparison, the continuous Floquet eigenvalue problem Eq.˜4 seeks eigenpairs $(u(0),e^{\mu T})$. The solution $x(t)$ of Eq.˜1 with $x(0)=u(0)$ satisfies $x(T)=e^{\mu T}x(0)$. Accordingly, one may associate the Floquet multipliers $e^{\mu T}$ with a subset of eigenvalues of $F_{p,0}$, in the sense that such eigenvalues are expected to approximate $e^{\mu T}$ as the grid is refined.
Importantly, the $n$ Floquet multipliers and their associated eigenvectors determine only $n$ reference eigenpairs of $F_{p,0}$. The remaining (d1)n(d-1)n eigenpairs do not admit a direct continuous-time interpretation and arise from the augmented state space induced by the $d$-step discretization.
4.2 Convergence Results of Periodic Eigenpairs of pPEP
The discussion in the previous subsection suggests that only $n$ eigenpairs of the one-period stacked propagator $F_{p,0}$ are relevant for approximating the Floquet multipliers of the underlying LTP system, while the remaining eigenvalues arise from the augmented state space induced by the multistep discretization. We now make this correspondence precise by establishing rigorous convergence results for the periodic eigenpairs of the pPEP.
To ensure that these convergence results are well-defined under time-grid refinement, we impose several standard assumptions on the multistep discretization.
  1. (A1)
    Smoothness of Multistep Method Coefficients. We assume that the functions $\alpha_{j}(\boldsymbol{\omega})$ and $\beta_{j}(\boldsymbol{\omega})$ are $C^{1}$-continuous with respect to $\boldsymbol{\omega}$ in a neighborhood of the uniform grid ratio $\mathbf{1}=(1,\ldots,1)$.
  2. (A2)
    Consistency of the Multistep Method. The $d$-step multistep method has consistency order $s\geq 1$. Thus, the coefficients satisfy $\sum_{j=0}^{d}\alpha_{j}^{(i)}=0$.
  3. (A3)
    Strong Stability of the Multistep Method. The underlying constant-stepsize method is strongly stable in the sense of [8]. Specifically, the first generating polynomial
    $\rho(\nu)=\sum_{j=0}^{d}\alpha_{j}(\mathbf{1})\nu^{j}$
    has a simple root at $\nu_{1}=1$, referred to as the principal root, while the remaining roots $\nu_{j}$, $j=2,\ldots,d$, referred to as the parasitic roots, satisfy $|\nu_{j}|<1$ [17]. Moreover, all parasitic roots are assumed to be distinct.
  4. (A4)
    Bounded Total Deviation of the Periodic Time Grid. The step ratios are uniformly bounded, more over, they satisfy $\sum_{i=0}^{p}\lvert\omega^{(i)}-1\rvert\leq C_{T},$ where $C_{T}>0$ is independent of the time grid refinement.
The assumptions (A1)–(A4) are standard in the classical theory of variable stepsize multistep methods for initial value problems [17]. Under these assumptions, the following theorem states the main convergence result for the periodic eigenpairs of the pPEP.
Theorem 4.1 (Convergence of Floquet Multipliers).
Assume that the periodic matrix $G(t)$ in Eq.˜1 is continuously differentiable on $[0,T]$, and hence uniformly bounded,
$\|G(t)\|_{2}\leq C_{G}.$
All Floquet multipliers $e^{\mu_{k}T}$, $k=1,\ldots,n$, are semisimple and ordered by non-increasing modulus. Suppose that the multistep method satisfies Assumptions (A1)(A4).
Then, as the maximum stepsize $h\to 0$ (equivalently, as $p\to\infty$), the spectrum of the one-period stacked propagator $F_{p,0}$ admits a decomposition into a principal part and a parasitic part with the following properties.
  1. 1.
    There exist $n$ eigenvalues of $F_{p,0}$, denoted by $\{\lambda_{k}(h)\}_{k=1}^{n}$, such that, after suitable ordering,
    $\lambda_{k}(h)=e^{\mu_{k}T}+O(h^{s}),\qquad k=1,\ldots,n,$
    where $\{e^{\mu_{k}T}\}_{k=1}^{n}$ are the Floquet multipliers of the underlying LTP system.
  2. 2.
    For any fixed $m$, $1\leq m\leq n$, the invariant subspace spanned by the eigenvectors associated with the $m$ dominant eigenvalues $\{\lambda_{k}(h)\}_{k=1}^{m}$, ordered by non-increasing modulus, converges to the subspace spanned by the first $m$ columns of $\mathcal{X}(t_{0})$ at the rate $O(h^{s}/\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert>0)$ if $\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert>0$.
  3. 3.
    The remaining (d1)n(d-1)n eigenvalues of $F_{p,0}$, referred to as parasitic periodic eigenvalues, satisfy
    $|\lambda_{k}(h)|\leq C\,\nu_{\max}^{p},\qquad k=n+1,\ldots,nd,\quad\nu_{\max}=\max_{2\leq j\leq d}|\nu_{j}|<1.$
    They therefore converge to 0 at a geometric rate as $p\to\infty$.
4.3 Construction of the Moving Basis and the Reference Transition Operator
To analyze the variable stepsize case, we construct a moving basis $\mathcal{Z}(t_{i})$ in the augmented state space $\mathbb{R}^{nd}$. Let $X(t)=\Phi(t,0)U(0)\in\mathbb{R}^{n\times n}$ represent the principal solution component, where $\Phi(t,0)$ is the state transition matrix of the ODE and $U(0)$ contains the Floquet eigenvectors. The parasitic components are modeled explicitly by the roots $\nu_{j}$. We define the moving basis matrix $\mathcal{Z}(t_{i})\in\mathbb{R}^{nd\times nd}$ by concatenating the snapshots of the principal solution and the parasitic modes:
(23) $\mathcal{Z}(t_{i})\equiv$ $\begin{bmatrix}X(t_{i-d+1})&\nu_{2}^{i-d+1}I_{n}&\cdots&\nu_{d}^{i-d+1}I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ X(t_{i-1})&\nu_{2}^{i-1}I_{n}&\cdots&\nu_{d}^{i-1}I_{n}\\ X(t_{i})&\nu_{2}^{i}I_{n}&\cdots&\nu_{d}^{i}I_{n}\end{bmatrix}$
$=$ $\begin{bmatrix}\mathcal{X}(t_{i})&\mathcal{Y}_{2}(t_{i})&\cdots&&\mathcal{Y}_{d}(t_{i}),\end{bmatrix}$
where $\mathcal{X}(t_{i})$ is the first block column of $\mathcal{Z}(t_{i})$ and $\mathcal{Y}_{\tau}$ denotes the $\tau$-th block column of $\mathcal{Z}(t_{i})$ that $\tau=2,\ldots,d$. The basis $\mathcal{Z}(t_{i})$ captures the state of the multistep method at step $i$, effectively separating the principal subspace from the parasitic subspaces. We observe that $\mathcal{Z}(t_{i})$ admits the decomposition:
(24) $\mathcal{Z}(t_{i})=\underbrace{\begin{bmatrix}\Phi(t_{i-d+1},t_{i})&I_{n}&\cdots&I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ \Phi(t_{i-1},t_{i})&\nu_{2}^{d-1}I_{n}&\cdots&\nu_{d}^{d-1}I_{n}\\ I_{n}&\nu_{2}^{d}I_{n}&\cdots&\nu_{d}^{d}I_{n}\end{bmatrix}}_{\mathcal{V}(t_{i})}\operatorname{diag}\left(X(t_{i}),\,\nu_{2}^{i-d}I_{n},\,\ldots,\,\nu_{d}^{i-d}I_{n}\right).$
As the stepsize approaches zero, $\Phi(t_{i-j},t_{i})$ approaches to $I_{n}$ with $\mathcal{O}(h)$. Consequently, the block matrix $\mathcal{V}(t_{i})$ converges to a block Vandermonde matrix. Since the roots $\{1,\nu_{2},\dots,\nu_{d}\}$ are distinct, $\mathcal{V}(t_{i})$ is nonsingular for sufficiently small stepsizes. This guarantees that $\mathcal{Z}(t_{i})$ is invertible.
Given the nonsingular bases at consecutive time steps, the corresponding reference state-transition operator $\Psi(t_{i},t_{i-1})$ is well defined.
(25) $\Psi(t_{i},t_{i-1})\equiv\mathcal{Z}(t_{i})\,[\mathcal{Z}(t_{i-1})]^{-1}.$
Noting that $X(t_{i+p})=X(t_{i})e^{MT}$, the moving basis $\mathcal{Z}(t_{i})$ is periodic up to a multiplying a diagonal matrix:
(26) $\mathcal{Z}(t_{i+p})=\mathcal{Z}(t_{i})\,\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix},\quad D=\operatorname{diag}\left(\nu_{2}^{p},\ldots,\nu_{d}^{p}\right)\otimes I_{n}.$
Thus, the reference state-transition operator over one period possesses the eigendecomposition:
(27) $\Psi(t_{i+p},t_{i})\,\mathcal{Z}(t_{i})=\mathcal{Z}(t_{i})\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}$.
4.4 Convergence Analysis
The construction of the reference stacked propagator $\Psi(t_{i+p},t_{i})$ establishes the framework for analyzing the discrete stacked propagator $F_{i+p,i}=L^{(i+p)}\cdots L^{(i+1)}$. Setting $i=0$ without loss of generality, the analysis proceeds by comparing $F_{p,0}$ against $\Psi(t_{p},t_{0})$ in a three-stage process:
  1. 1.
    Local Estimation: Compute the local deviation between the local discrete transition $L^{(i)}$ and its reference $\Psi(t_{i},t_{i-1})$ on the moving basis:
    (28) $\left[\mathcal{Z}(t_{i})\right]^{-1}L^{(i)}\mathcal{Z}(t_{i-1})=I_{dn}+\Delta_{i}.$
  2. 2.
    Global Accumulation: Compute the global deviation between the one-period discrete propagator $F_{p,0}=L^{(p)}\cdots L^{(1)}$ and its reference $\Psi(t_{p},t_{0})$ on the moving basis:
    $\mathcal{Z}(t_{0})]^{-1}F_{p,0}\mathcal{Z}(t_{0})=\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}(I_{dn}+\Delta),$
    where $\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}$ is the eigenvalues of $\Psi(t_{p},t_{0})$ and $\Delta$ represents the accumulated global perturbation.
  3. 3.
    Eigenpair Perturbation Estimation: Since the representation of $\Psi(t_{p},t_{0})$ on the moving basis is an eigendecomposition Eq.˜27, then apply the subspace and eigenvalue perturbation theory to estimate how eigenpair of $F_{p,0}$ deviates from $\Psi(t_{p},t_{0})$.
Lemma 4.2 (Local Difference Estimation).
Suppose that Assumptions (A1)–(A4) hold. Then there exists a threshold $\delta>0$ such that, for any periodic grid satisfying $\max_{i}h^{(i)}\leq\delta$ the local deviation of $L^{(i)}$ from $\Psi(t_{i},t_{i-1})$ admits the representation
$\Delta_{i}\equiv\mathcal{Z}^{-1}(t_{i})\bigl[L^{(i)}-\Psi(t_{i},t_{i-1})\bigr]\mathcal{Z}(t_{i-1})=\begin{bmatrix}E_{i,\mathcal{X}\mathcal{X}}&E_{i,\mathcal{X}\mathcal{Y}}\\ E_{i,\mathcal{Y}\mathcal{X}}&E_{i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}$,
where the block partition is consistent with the decomposition of $\mathcal{Z}$ into its principal and parasitic components, with $E_{i,\mathcal{X}\mathcal{X}}\in\mathbb{R}^{n\times n}$ and $E_{i,\mathcal{Y}\mathcal{Y}}\in\mathbb{R}^{(d-1)n\times(d-1)n}.$
Moreover, there exist constants $C_{1},C_{2},K>0$, independent of $i$, such that
$\left\|\begin{bmatrix}E_{i,\mathcal{X}\mathcal{X}}\\ E_{i,\mathcal{X}\mathcal{Y}}\end{bmatrix}\right\|_{2}\leq C_{1}\,{h^{(i)}}^{s+1},\qquad\left\|\begin{bmatrix}E_{i,\mathcal{Y}\mathcal{X}}\\ E_{i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\|_{2}\leq C_{2}\,h^{(i)}+K\sum_{l=0}^{d}\bigl\lvert\omega^{(i-d+l)}-1\bigr\rvert$.
Proof 4.3.
By Assumption (A4), the step ratios $\omega^{(i)}$ remain uniformly bounded. Together with the smooth dependence of the coefficients $\alpha_{j}(\boldsymbol{\omega})$ and $\beta_{j}(\boldsymbol{\omega})$ Assumption (A1), this implies uniform boundedness of $\alpha_{j}^{(i)}$ and $\beta_{j}^{(i)}$. Hence, there exists $\delta>0$ sufficiently small such that for $h^{(i)}<\delta$, the matrix $\alpha_{d}^{(i)}I_{n}-h^{(i)}\beta_{d}^{(i)}G^{(i)}$ is invertible with uniformly bounded inverse.
We first examine the action of $L^{(i)}-\Psi(t_{i},t_{i-1})$ on each block column of $\mathcal{Z}(t_{i-1})$.
$\left[L^{(i)}-\Psi(t_{i},t_{i-1})\right]\mathcal{Z}(t_{i-1})$
$=$ $\left[L^{(i)}-\Psi(t_{i},t_{i-1})\right]\left[\mathcal{X}_{1}(t_{i-1}),\,\mathcal{Y}_{2}(t_{i-1}),\,\ldots,\,\mathcal{Y}_{d}(t_{i-1})\right]$
$=$ $\begin{bmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ R_{\mathcal{X}_{i}}&R_{\mathcal{Y}_{2,i}}\nu_{2}^{i-d}&\cdots&R_{\mathcal{Y}_{d,i}}\nu_{d}^{i-d}\end{bmatrix}$.
By Assumption (A2) and the uniform boundedness of the inverse matrix,
$\left\lVert R_{\mathcal{X}_{i}}\right\rVert_{2}\leq C_{1}\,{h^{(i)}}^{s+1}.$
For $\tau=2,\ldots,d$, the residual takes the form
$R_{\mathcal{Y}_{\tau,i}}=-\left(\alpha_{d}^{(i)}I_{n}-h^{(i)}\beta_{d}^{(i)}G^{(i)}\right)^{-1}\left(\sum_{j=0}^{d}h^{(i)}\beta_{j}^{(i)}G^{(i-d+j)}\nu_{\tau}^{j}+\sum_{j=0}^{d}\alpha_{j}^{(i)}\nu_{\tau}^{j}I_{n}\right).$
The first term inside the right bracket is $O(h^{(i)})$ by boundedness of the coefficients $\beta_{j}^{(i)}$ and $G(t)$. For the second term, using the parasitic root condition $\sum_{j}\alpha_{j}(\mathbf{1})\nu_{\tau}^{j}=0$ and the smoothness assumption on the coefficients (A1), we obtain
$\left\lvert\sum_{j=0}^{d}\alpha_{j}^{(i)}\nu_{\tau}^{j}\right\rvert=\left\lvert\sum_{j=0}^{d}\left(\alpha_{j}^{(i)}-\alpha_{j}(\boldsymbol{1})\right)\nu_{\tau}^{j}\right\rvert\leq K\sum_{l=0}^{d}|\omega^{(i-d+l)}-1|.$
Consequently,
$\left\lVert R_{\mathcal{Y}_{\tau,i}}\right\rVert_{2}\leq C_{2}h^{(i)}+K\sum_{l=0}^{d}|\omega^{(i-d+l)}-1|.$
We now transform the residual matrix to the moving basis by left-multiplying with $\mathcal{Z}^{-1}(t_{i})$. Therefore,
$\Delta_{i}$ $=\mathcal{Z}^{-1}(t_{i})\left[L^{(i)}\mathcal{Z}(t_{i-1})-\mathcal{Z}(t_{i})\right]$
$=\begin{bmatrix}X(t_{i-d+1})&I_{n}&\cdots&I_{n}\\ \vdots&\vdots&\ddots&\vdots\\ X(t_{i-1})&I_{n}\nu_{2}^{d-1}&\cdots&I_{n}\nu_{d}^{d-1}\\ X(t_{i})&I_{n}\nu_{2}^{d}&\cdots&I_{n}\nu_{d}^{d}\end{bmatrix}^{-1}\begin{bmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ R_{\mathcal{X}_{i}}&R_{\mathcal{Y}_{2,i}}&\cdots&R_{\mathcal{Y}_{d,i}}\end{bmatrix}$
inherits the same block-wise bounds as the residual matrix.
Finally, transforming to the $\mathcal{Z}$-coordinates via $\Delta_{i}=\mathcal{Z}^{-1}(t_{i})R_{i}$ and using the uniform boundedness of $\mathcal{Z}^{-1}(t_{i})$ yields the stated block-wise estimates.
Lemma 4.4 (Global Difference Estimation).
Suppose that Assumptions (A1)–(A4) hold. Then, with respect to the moving basis $\mathcal{Z}(t_{0})$, the discrete stacked propagator over one period $F_{p,0}$ admits the representation
[…]
where the block structure is consistent with the decomposition of $\mathcal{Z}(t_{0})$ into principal and parasitic components.
Moreover, under Assumptions (A1)–(A4), there exist constants $C_{1},C_{2}>0$, independent of the periodic time grid, such that
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{X}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{1}h^{s},\qquad\left\lVert\begin{bmatrix}E_{\mathcal{Y}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq C_{2}.$
Proof 4.5.
We begin by expressing $F_{p,0}$ in the basis $\mathcal{Z}(t_{0})$.
$\begin{aligned} &\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left[\mathcal{Z}^{-1}(t_{p})L^{(p)}\mathcal{Z}(t_{p-1})\right]\cdots\left[\mathcal{Z}^{-1}(t_{1})L^{(1)}\mathcal{Z}(t_{0})\right]\\ =&\mathcal{Z}^{-1}(t_{0})\mathcal{Z}(t_{p})\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\\ =&\begin{bmatrix}e^{MT}&\\ &D^{p}\end{bmatrix}\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)\end{aligned}.$
We then define the global error matrix and apply a telescoping sum expansion:
$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}&E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{X}}&E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\equiv$ $\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{1}\right)-I_{dn}$
$=$ $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i+1})\Delta_{p-i}.$
We now estimate the two block columns separately. For the first block column, the telescoping representation yields
$\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}$= $\sum_{i=0}^{p-1}(I_{dn}+\Delta_{p})\cdots(I_{dn}+\Delta_{p-i})\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}$.
By Lemma˜4.2, there exists a constant $C>0$ such that
$\left\lVert\left(I_{dn}+\Delta_{p}\right)\cdots\left(I_{dn}+\Delta_{p-i}\right)\right\rVert_{2}\leq\prod_{k=p-i}^{p}\left\lVert I_{dn}+\Delta_{k}\right\rVert_{2}\leq\prod_{k=p-i}^{p}(1+Ch^{(k)})\leq e^{CT},$
which is uniformly bounded for all $i=1,\ldots,p$.
Therefore, we obtain the estimate
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{X}}\\ E_{\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{X}}\\ E_{p-i,\mathcal{Y}\mathcal{X}}\end{bmatrix}\right\rVert_{2}\leq e^{CT}\sum_{i=0}^{p-1}(h^{(p-k)})^{s+1}\leq e^{CT}h^{s}$
where we used $\sum_{i=1}^{p}h^{(i)}=T$ and $h=\max_{i}h^{(i)}$.
An analogous argument applies to the second block column,
$\left\lVert\begin{bmatrix}E_{\mathcal{X}\mathcal{Y}}\\ E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}\leq$ $\sum_{i=1}^{p}e^{CT}\left\lVert\begin{bmatrix}E_{p-i,\mathcal{X}\mathcal{Y}}\\ E_{p-i,\mathcal{Y}\mathcal{Y}}\end{bmatrix}\right\rVert_{2}$
[…]
By Assumption (A4), the cumulative deviation of the step ratios is uniformly bounded, and therefore the right-hand side is bounded by a constant. This completes the proof.
The global difference estimates established above quantify the deviation between the one-period discrete propagator $F_{p,0}$ and its reference counterpart constructed $\Psi(t_{p},t_{0})$. To link it to estimates of eigenpairs difference, we recall classical results from invariant subspace perturbation theorem [22, 3].
Theorem 4.6 (Invariant Subspace Perturbation Theorem [22]).
Let $A,E\in\mathbb{C}^{n\times n}$, $P=[P_{1},P_{2}]$ is a unitary matrix. $P_{1}\in\mathbb{C}^{n\times l},(1\leq l\leq n-1)$. Assume $\operatorname{span}\left\{P_{1}\right\}$ is the invariant subspace of $A$. Let $P^{\mathrm{H}}AP$ and $P^{\mathrm{H}}EP$ are consistently separated as
$P^{\mathrm{H}}AP=\begin{bmatrix}A_{11}&A_{12}\\ &A_{22}\end{bmatrix},\quad P^{\mathrm{H}}EP=\begin{bmatrix}E_{11}&E_{12}\\ E_{21}&E_{22}\end{bmatrix}$
If $\delta\equiv\operatorname{sep}(A_{11},A_{22})-\left(\left\lVert E_{11}\right\rVert+\left\lVert E_{22}\right\rVert\right)>0$, and $\left\lVert E_{21}\right\rVert\left(\left\lVert A_{12}\right\rVert+\left\lVert E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}$, then there exist a matrix $R\in\mathbb{C}^{(n-l)\times l}$ satisfying $\left\lVert R\right\rVert\leq 2\delta^{-1}\left\lVert E_{21}\right\rVert$, such that $\operatorname{span}\left\{P\begin{bmatrix}I\\ R\end{bmatrix}\right\}$ is an invariant subspace of $A+E$.
Remark 4.7.
The separation between two square matrices $A_{11}$ and $A_{22}$ is defined as
$\operatorname{sep}_{2}(A_{11},A_{22}):=\inf_{\|X\|_{2}=1}\|A_{11}X-XA_{22}\|_{2}.$
If $A_{11}$ and $A_{22}$ are both normal, then the separation coincides with the spectral gap between their spectra.
$\operatorname{sep}_{2}(A_{11},A_{22})=\mathrm{d}(A_{11},A_{12})\equiv\min\{\left\lvert\lambda-\mu\right\rvert,\,\lambda\in\lambda(A_{11}),\mu\in\lambda(A_{22})\}.$
With the global difference estimate and the invariant subspace perturbation theorem at hand, we are now in a position to complete the proof of Theorem˜4.1.
Proof 4.8.
The three statements are proved by the same perturbation argument based on the block representation of $\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$ and a spectral separation estimate. We therefore present the proof for part (3), which contains the main ideas. Parts (2) and (1) follow by the same reasoning with minor simplifications, as outlined at the end of the proof.
Let $1\leq k\leq n$, we partition $\mathcal{X}(t_{0})$ as $\mathcal{X}(t_{0})=\begin{bmatrix}X_{k}(t_{0})&X_{n-k}(t_{0})\end{bmatrix}$ After the similarity transformation induced by the moving basis $\mathcal{Z}(t_{0})$, the one-period stacked propagator admits the block representation Eq.˜29:
$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})$
$=$ $\begin{bmatrix}e^{M_{k}T}&&\\ &e^{M_{n-k}T}\\ &&D^{p}\end{bmatrix}+\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\\ e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}&D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
For brevity, we set
$E_{11}=e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}},\quad E_{12}=\begin{bmatrix}e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{n-k}}&e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{Y}}\end{bmatrix}$,
$E_{21}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{k}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{k}}\end{bmatrix},\quad E_{22}=\begin{bmatrix}e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{X}_{n-k}}&e^{M_{n-k}T}E_{\mathcal{X}_{n-k}\mathcal{Y}}\\ D^{p}E_{\mathcal{Y}\mathcal{X}_{n-k}}&D^{p}E_{\mathcal{Y}\mathcal{Y}}\end{bmatrix}$.
Since all Floquet multipliers $e^{\mu_{k}T}$ are assumed semisimple and sorted by non-increasing order, thus $e^{\mu_{n}T}$ is the smallest magnitude Floquet multiplier. Since $\|D^{p}\|=\nu_{\max}^{p}\to 0$ as $p\to\infty$, there exists $p_{0}\in\mathbb{N}$ such that $\|D^{p}\|\leq e^{\mu_{n}T}/2$ for all $p\geq p_{0}$. Because both $e^{MT}$ and $D^{p}$ are diagonal (hence normal), their spectra are separated, and
$\operatorname{sep}(e^{MT},D^{p})=\operatorname{dist}\!\bigl(\lambda(e^{MT}),\lambda(D^{p})\bigr)\geq e^{\mu_{n}T}/2,\qquad p\geq p_{0}.$
For sufficient large $p$, we have
$\begin{aligned} \delta=&\operatorname{sep}(e^{M_{k}T},\begin{bmatrix}e^{M_{n-k}}&\\ &D^{p}\end{bmatrix})-\left(\left\lVert e^{M_{k}T}E_{\mathcal{X}_{k}\mathcal{X}_{k}}\right\rVert+\left\lVert\begin{bmatrix}E_{11}&\\ &E_{22}\end{bmatrix}\right\rVert\right)\\ >&1/2\left\lvert e^{\mu_{m}T}-e^{\mu_{m+1}T}\right\rvert\end{aligned}$,
and
$\left\lVert E_{21}\right\rVert\left(\left\lVert 0\right\rVert+\left\lVert E_{21}+E_{12}\right\rVert\right)<\frac{1}{4}\delta^{2}.$
Thus, the conditions of Theorem˜4.6 are satisfied, and there exists a matrix $R\in\mathbb{C}^{((d-1)n+k)\times k}$ with
$\|R\|\leq C\,\nu_{\max}^{p},$
such that
$\mathcal{Z}^{-1}(t_{0})F_{p,0}\mathcal{Z}(t_{0})\begin{bmatrix}I_{k}\\ R\end{bmatrix}=\begin{bmatrix}I_{k}\\ R\end{bmatrix}B,$
where the reduced $n\times n$ matrix $B$ is given by
$B:=e^{M_{k}T}+e^{M_{k}T}E_{11}-e^{M_{k}T}E_{12}R.$
Thus,
$F_{p,0}\left(\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R\right)=\left(\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R\right)B,$
that is $\mathcal{X}_{k,0}=\mathcal{X}_{k}(t_{0})+\begin{bmatrix}\mathcal{X}_{n-k}(t_{0})&\mathcal{Y}(t_{0})\end{bmatrix}R$, thus we proved (3).
Using $\|E_{11}\|=O(h^{s})$, $\|E_{12}\|=O(h^{s})$ and $\|R\|=O(\nu_{\max}^{p})$, we obtain
$\|B-e^{M_{k}T}\|\leq Ch^{s}+C\nu_{\max}^{p}.$
Since $e^{M_{k}T}$ is diagonal (hence normal), its eigenvalues are stable under perturbations:
$\operatorname{dist}\!\bigl(\lambda(B),\lambda(e^{M_{k}T})\bigr)\leq\|B-e^{M_{k}T}\|.$
Therefore, after suitable ordering, the $n$ eigenvalues of $F_{p,0}$ associated with the dominant spectral cluster satisfy
$\lambda_{k}(h)=e^{\mu_{k}T}+O(h^{s})+O(\nu_{\max}^{p}),\qquad k=1,\ldots,n.$
Because $\nu_{\max}^{p}$ decays geometrically as $p\to\infty$, this yields the claimed $O(h^{s})$ convergence of the principal eigenvalues. Let $k=n$, we proved $(1)$.
The remaining (d1)n(d-1)n eigenvalues correspond to the complementary block and arise from a perturbation of $D^{p}$ of size $O(\nu_{\max}^{p})$. Hence each parasitic eigenvalue satisfies
$|\lambda_{k}|\leq C\|D^{p}\|\leq C\nu_{\max}^{p},\quad k=n+1,\ldots,dn,$
which proves the geometric decay of the parasitic spectrum.
5 Algorithm and Complexity Analysis
We present pTOAR, an iterative method for computing the dominant $k$ eigenpairs of the pPEP Eq.˜14. It combines the periodic Arnoldi process [18] with TOAR style basis compression [19] to reduce memory and computational cost, making it suitable for large-scale problems. The remainder of this section details the algorithmic steps and its underlying principles.
The periodic Krylov-Schur (pKS) algorithm [18] is well-suited for the linearized pPEP Eq.˜19. It avoids explicitly forming the full product $F_{i+p,i}=L^{(p)}\cdots\allowbreak L^{(1)}$, while maintaining accuracy even for eigenvalues spanning several orders of magnitude. The pKS algorithm constructs a periodic Arnoldi process:
(30) $L^{(i)}V^{(i-1)}=V^{(i)}T^{(i)},$
where $V^{(i)}$ for $i=1,\ldots,p-1$ are $(nd)\times k$ orthonormal matrices, and $V^{(p)}\in\mathbb{C}^{nd\times(k+1)}$, which includes the additional Arnoldi residual vector. The projected matrices $T^{(i)}\in\mathbb{C}^{k\times k}$ are upper triangular, except for upper-Hessenberg matrix $T^{(p)}\in\mathbb{C}^{(k+1)\times k}$. Although the basis dimensions slightly differ at the boundaries, such variations do not affect the asymptotic memory complexity. For simplicity, we omit these boundary details in the subsequent discussion.
A limitation of pKS is that it disregards the linearized structure of $L^{(i)}$, resulting in a dimension increase from $n$ to $nd$ and leading to unnecessary computational overhead. To address this issue, we draw inspiration from the Two-level Orthogonal Arnoldi (TOAR) method [20], which was originally developed for polynomial eigenvalue problems to compress the basis arising from linearization.
We introduce pTOAR, a hybrid of the pKS method and TOAR-style compression. The key difference is that pTOAR stores the periodic Arnoldi basis $V^{(i)}$ in pKS in an implicit, compact form, greatly saving memory. We now show the details of the compression mechanism.
We explicitly substitute the block form of $L^{(i)}$ into Eq.˜30, and partition the $V^{(i)}$ following the block structure of $L^{(i)}$, as shown in Eq.˜31:
(31) $\begin{bmatrix}&I_{n}&&\\ &&\ddots&\\ &&&I_{n}\\ A_{0}^{(i)}&A_{1}^{(i)}&\cdots&A_{d-1}^{(i)}\\ \end{bmatrix}\begin{bmatrix}V_{1}^{(i-1)}\\ \vdots\\ V_{d-1}^{(i-1)}\\ V_{d}^{(i-1)}\end{bmatrix}=\begin{bmatrix}V_{1}^{(i)}\\ \vdots\\ V_{d-1}^{(i)}\\ V_{d}^{(i)}\end{bmatrix}T^{(i)}.$
The key observation of pTOAR from Eq.˜31 is an information overlap relationship across adjacent basis matrices:
$V_{j+1}^{(i)}T^{(i)}=V_{j}^{(i+1)},\,j=1,\ldots,d-1,\,i=1,\ldots,p.$
This implies that a single subspace can represent $d$ block segments drawn from $d$ consecutive basis matrices. Specifically, we define a shared basis $Q^{(i)}\in\mathbb{C}^{n\times k}$ such that
$\operatorname{span}(Q^{(i)})=\operatorname{span}(V_{d}^{(i)},V_{d-1}^{(i+1)},\ldots,V_{1}^{(i+d-1)}).$
Utilizing this structure, each Arnoldi basis $V^{(i)}$ can be compactly expressed by form:
$V^{(i)}=\begin{bmatrix}Q^{(i-d+1)}&&\\ &\ddots&\\ &&Q^{(i)}\end{bmatrix}\begin{bmatrix}U_{1}^{(i)}\\ \vdots\\ U_{d}^{(i)}\end{bmatrix}$,
where each $\left\{Q^{(j)}\right\}_{j=i-d+1}^{i}\in\mathbb{C}^{n\times k}$ is an orthonormal basis representing the shared subspace, and $\left\{U_{j}^{(i)}\right\}_{j=1}^{d}\in\mathbb{C}^{k\times k}$ are local representation matrices. It follows that the block-diagonal matrix $\operatorname{blkdiag}(Q^{(i-d+1)},\dots,Q^{(i)})$ is also orthogonal, which in turn implies that the tall matrix $\begin{bmatrix}(U_{1}^{(i)})^{\mathrm{T}}&\dots&(U_{d}^{(i)})^{\mathrm{T}}\end{bmatrix}^{\mathrm{T}}$ is also orthonormal.
The pTOAR algorithm proceeds through periodic Arnoldi basis construction, Ritz value extraction and reordering, and optional restarts, following a workflow analogous to that of pKS but differing in implementation details. We will discuss these differences step by step, and then estimate the memory and computational costs for both algorithms. Because a restart simply repeats this core procedure and its occurrence depends on the eigenvalue problem, and most practical cases requires no or only once restart, we estimate the costs by examining the non-restarted case.
Similar to pKS, the periodic Arnoldi basis update in pTOAR involves a matrix-vector product $L^{(i)}v^{(i-1)}$, whose result is used to update $Q^{(i)}$ and $U_{\cdot,j}^{(i)}$. Owing to the linearization structure of $L^{(i)}$, each product updates only the last block:
$L^{(i)}v^{(i-1)}=\left[{v_{2}^{(i-1)}}^{\mathrm{H}},\cdots,{v_{d}^{(i-1)}}^{\mathrm{H}},{v_{d}^{(i)}}^{\mathrm{H}}\right]^{\mathrm{H}},\,v_{d}^{(i)}={A_{d}^{(i)}}^{-1}\left(\sum_{j=0}^{d-1}A_{j}^{(i)}v_{j}^{(i-1)}\right),$
where $v^{(i-1)}$ is the last column of $V^{(i-1)}$, and $v_{j}^{(i-1)}$ is its $j$-th block of $v^{(i-1)}$, also the last column of $V_{j}^{(i-1)}$ in Eq.˜31.
The second step in periodic Arnoldi basis construction is orthogonalization. Once $v_{d}^{(i)}$ is obtained, the pKS constructs and orthogonalizes $v^{(i)}$ against the current full basis $V^{(i)}$. In contrast, pTOAR first orthogonalizes $v_{d}^{(i)}$ against the current basis $Q^{(i)}$, and augments $Q^{(i)}$ with the normalized component. The representation of $v_{d}^{(i)}$ on the updated $Q^{(i)}$ is appended to the right of $U_{d}^{(i)}$. The remaining $U_{j}^{(i)}$ are extended from shifted representations $v_{j+1}^{(i-1)}$ in $Q^{(i-d+j)}$.
The computation of $v_{d}^{(i)}$ requires $d$ matrix-vector products, and one linear system solve, all in dimension $n$. Assuming matrix-vector products cost as $\mathcal{O}(n^{\gamma})$ and linear system solves cost as $\mathcal{O}(n^{\gamma+1})$, the total cost of constructing the periodic Arnoldi basis over $k$ iterations is $\mathcal{O}\left(pk(dn^{\gamma}+n^{\gamma+1})\right)$, where $\gamma=1$ for sparse and $\gamma=2$ for dense systems. In pTOAR, each $v_{j}^{(i-1)}$ needs to be explicitly reconstructed through $v_{j}^{(i-1)}=Q^{(i-d+j)}u_{j}^{(i-1)}$ which introduces an extra $2ndk$ cost per iteration. For orthogonalization, pKS applies classical Gram-Schmidt process to an $nd$-dimensional basis, incurring a cost of $2pndk^{2}+\mathcal{O}(pndk)$. In contrast, pTOAR orthogonalizes only the compressed basis $Q^{(i)}$, reducing this cost to $2pnk^{2}+\mathcal{O}(pnk)$. However, updating $U_{j}^{(i)}$ matrices introduces another $2pdk^{2}+\mathcal{O}(pdk)$ cost.
As for memory usage, instead of storing the all $p$ full basis $V^{(i)}\in\mathbb{C}^{nd\times k}$, pTOAR only stores $p$ subspace bases $Q^{(i)}$ and $pd$ representations $U_{j}^{(i)}$, reducing the overall memory cost from pndkpndk to $pnk+pdk^{2}$.
The extraction and reordering of Ritz values are identical for pKS and pTOAR. For a Krylov subspace of dimension $k$, this step applies the periodic QR algorithm, a direct method [5], to the projected periodic matrices $T_{k}^{(i)}$, incurring a computational cost of $\mathcal{O}(pk^{3})$ [20]. Each small projected matrix $T_{k}^{(i)}$ requires an additional $k^{2}$ memory, leading to a total of $pk^{2}$ across all $p$ slices.
At last, we evaluate memory requirement for storing the input data. For each $L^{(i)}$, it suffices to store the sampled matrices $G^{(i)}$ instead of the coefficient matrix $A_{j}^{(i)}$ in Eq.˜12 and the full $nd\times nd$ matrix $L^{(i)}$. This leads to a memory cost of $n^{\gamma}$ per slice, where $\gamma=1$ if the matrices $A_{j}^{(i)}$ are sparse, and $\gamma=2$ if dense. Combining this with the basis memory and the Ritz extraction memory, the overall memory cost is thus $pn^{\gamma}+pnkd$ under assumption $k\ll n$.
Table˜1 compares the computational and memory costs of three algorithmic setups: one-step collocation with pKS, and multistep methods with pKS or pTOAR. Each leads to a distinct periodic eigenvalue problem for computing the dominant $k$ Floquet multipliers. Costs are measured per full Arnoldi process, as convergence mainly depends on the Floquet multipliers’ spectral distribution rather than the solver details. Complexity estimates for the condensation step in the one-step collocation method are from [20].
Algorithm Collocation Multistep Multistep
+pKS +pKS +pTOAR
memory Cost
Coefficient Matrices $pn^{2}$ $pn^{\gamma}$ $pn^{\gamma}$
Arnoldi Basis kpnkpn kdpnkdpn kpnkpn
Total Memory $pn^{2}$ $kdpn+pn^{\gamma}$ $kpn+pn^{\gamma}$
Computational Cost
[…]
Arnoldi Process $\mathcal{O}(pn^{3}k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$
Orthogonalization $2k^{2}pn$ $2k^{2}dpn$ $2k^{2}p(n+d)$
Ritz Value Extraction $4pk^{3}$ $4pk^{3}$ $4pk^{3}$
Total FLOPs $\mathcal{O}(pn^{3}(d^{3}+k))$ $\mathcal{O}(pn^{\gamma}(n+d)k)$ $\mathcal{O}(pn^{\gamma}(n+d)k)$
Table 1: Comparison of memory and computational costs for different methods, where $\gamma=1$ if input matrices are sparse, and $\gamma=2$ if they are dense.
As shown in Table˜1, the collocation method suffers from the expensive condensation process and dense system matrices, while pKS and pTOAR avoid both. Compared to pKS, pTOAR further reduces memory and orthogonalization costs when $d>1$.
To compare with backward Euler, we can set $d=0$, since there is no extra matrix-vector product or memory cost in this case. For $d>1$, multistep methods incur extra matrix-vector products and memory for $pd$ matrices $U_{j}^{(i)}$, but this overhead is typically negligible compared to solving linear systems.
Finally, we are prepared to give the full algorithm for computing the Floquet multipliers via Discretization and iterative eigenvalue solvers.
Algorithm 1 Algorithm for Computing Dominant Floquet Eigenpairs with Multistep Discretization and Periodic Iterative Solvers
0:Time samples $\{t_{i}\}_{i=0}^{p}$ with $0=t_{0}<t_{1}<\cdots<t_{p}=T$, where $T$ is the period;Discrete evaluations $\{G(t_{i})\}_{i=0}^{p}$ of the system matrix; Target number of eigenpairs $k$.
0:Dominant $k$ Floquet multipliers and their corresponding eigenvectors.
1:Discretization: At each time $t_{i}$, define $\left\{A_{j}^{(i)}\right\}_{j=0}^{d}$ in Eq.˜12 implicitly via matrix-vector products routines, and invoke them to define matrix-vector products of $L^{(i)}$, as inEq.˜18.
2:Eigenpair Computation: Apply the pTOAR to compute the dominant $k$ eigenpairs of the linearized periodic EVP Eq.˜19.
6 Numerical Results
This section presents three numerical experiments. The first is an artificial example constructed from known Floquet eigenpairs to verify convergence theory and the distribution of parasitic eigenvalues.
The second and third experiments address two common scenarios where Floquet multipliers are needed. In the first scenario, the matrix $G(t;\delta)$ is given explicitly, and tracking the evolution of multipliers with respect to $\delta$ helps detect bifurcations or instability onset [20]. We test this scenario using a benchmark problem and compare our pTOAR method with existing packages MATCONT and AUTO-07p. In the second scenario, $G(t)$ is only available at discrete time samples, without an analytic form; here the focus is often on computing the Floquet eigenvector associated with the dominant multiplier (e.g., in circuit oscillator design) [9, 10]. We illustrate this scenario with an industrial circuit application, where fixed sampling makes collocation infeasible and highlights the advantage of our multistep-based approach.
All numerical experiments are conducted using MATLAB 2024b on a system equipped with an Intel Core i5-1135G7 CPU at 2.4 GHz and 16 GB of RAM.
We assess the accuracy of the computed Floquet multipliers $\lambda$ and eigenvectors $x$ using relative errors:
(32) $e_{\text{val}}=\frac{\left|\lambda-e^{\mu T}\right|}{\left|e^{\mu T}\right|},\quad e_{\text{vec}}=\frac{\|x_{p}-x(T)\|_{2}}{\|x(T)\|_{2}},$
where $e^{\mu T}$ and $x(T)$ are the exact Floquet multiplier and eigenvector. For the invariant subspace spanned by several eigenvectors, we define the subspace error between computed subspace $U$ and exact subspace $V$ by
(33) $e_{\text{subspace}}=\Theta(U,V),$
where $\Theta(U,V)$ is the principal angle between the two subspaces. When exact solutions are not available, we use higher-order multistep methods to generate reference solutions, whose discrepancy from the exact operator is negligible for our evaluation purposes.
6.1 Convergence Behavior of Floquet multipliers and Parasitic Eiegnvalues
Our first experiment is based on a two-dimensional LTP system given by Eq.˜34,
(34) $\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}\alpha\cos^{2}t+(\beta-2)\sin^{2}t&1+(\beta-\alpha-2)\sin t\cos t\\ -1+(\beta-\alpha-2)\sin t\cos t&\alpha\sin^{2}t+(\beta-2)\cos^{2}t\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$.
This model is obtained by linearizing the Stuart-Landau oscillator along its limit cycle. Its Floquet multipliers are $e^{2\pi\alpha}$ and $e^{2\pi(\beta-2)}$, and their corresponding eigenvectors are columns of
(35) $U(t)=\begin{bmatrix}\cos t&\sin t\\ -\sin t&\cos t\end{bmatrix}$.
We discretize one period $0,2\pi]$ with variable stepsize, and apply Gear2, Gear3 methods to obtain the periodic-EVP Eq.˜19. The Floquet multiplier, its associated eigenvector, and the parasitic eigenvalues from multistep methods are then computed using the periodic QR algorithm [5] with full precision. For reference, we also apply the backward Euler (BE) discretization, which does not produce spurio us eigenvalues, to serve as a baseline for convergence comparison.
Through this experiment, we set $\alpha=0.1,\beta=0.1$. The system is discretized over one period into $2^{k}$ points with stepsizes alternating in the ratio $1:2:1:2:$... We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize. Through this experiment, we set $\alpha=0.1$ and $\beta=0.1$. The system is discretized over one period into $2^{k}$ grid points with variable stepsizes. To ensure that the resulting time grid satisfies Assumption (A4), we evenly insert four steps where the stepsize is doubled and four steps where the stepsize is halved, so that the total deviation of the step ratios from unity remains bounded over one period. We study the convergence of the computed Floquet eigenpairs with respect to the largest stepsize.
(a) Floquet Eigenvalue Error vs. Maximum Discrete Stepsize
(b) Floquet Eigenvector Error vs. Maximum Discrete Stepsize
Figure 1: Convergence Behavior of Dominant Floquet Eigenpair
As shown in Fig.˜1(a) and Fig.˜1(b), BE, Gear2, and Gear3 exhibit first, second, and third order convergence, respectively, consistent with the theoretical prediction. Only the errors for the dominant Floquet multiplier and its eigenvector are shown here, as the results for the other mode are similar.
(a) Parasitic Eigenvalues of Gear2
(b) Parasitic Eigenvalues of Gear3
Figure 2: Convergence Behavior of Parasitic Eigenvalues
Figures˜2(a) and 2(b) illustrate the convergence of the parasitic eigenvalues of Gear2 (two eigenvalues) and Gear3 (four eigenvalues), respectively, as the stepsize decreases. Since these eigenvalues are very small, their logarithms are shown, with different markers indicating individual modes. The red lines represent $\log(\nu)$ and serve as a reference for the expected decay rate. As each parasitic eigenvalue equals a Floquet multiplier of moderate magnitude times $\nu^{p}$, we compare only orders of magnitude rather than exact values.
The figures show that the computed parasitic eigenvalues match the expected magnitude very well, far exceeding the prediction of Theorem˜4.1, which guarantees only linear decay with $h$.
6.2 Scenario 1: Coupled Stuart-Landau Oscillators
The first numerical experiment uses a coupled Stuart-Landau system adapted from [13, 20], whose analytical form is given by
[…]
We fix $\beta=0.5$ and use $\delta$ as the continuation parameter, approaching $\delta=0.25$ from below. Our focus is on the Floquet multipliers of the variational system associated to a limit cycle in Eq.˜36, where two oscillators are $\pi$ radians out of phase, i.e., $x_{1}=-x_{2},y_{1}=-y_{2}$. As $\delta$ approaches $0.25$, this limit cycle approaches a heteroclinic loop, and the period grows raplidly, and the Floquet multipliers vary sharply. Our goal is to compute its Floquet multipliers with high accuracy to determine their position relative to the unit circle in the complex plane.
For a given pair of $(\delta,\beta)$, the periodic solution is first computed using existing continuation software—MATCONT [11] in MATLAB and AUTO-07p [12] in FORTRAN. These tools provide not only the periodic orbit but also the Jacobian along the trajectory, enabling us to study the associated variational system.
We test three discretization and eigensolver combinations on the coupled Stuart-Landau system Eq.˜36, as listed in Table˜2.
Table 2: Comparison of Software, Discretization Methods, and Algorithms
Software Discretization Method Algorithm
AUTO-07p Collocation O.K.Floquet [13]
Modified MATCONT Collocation Periodic QR[5]
pTOAR Multistep Method pTOAR
The first method uses the built-in Floquet multiplier computation in AUTO-07p, which applies a collocation discretization and the O.K. Floquet algorithm [13]. The second method modifies MATCONT: it retains the collocation discretization but replaces the built-in MATLAB eig eigenvalue solver with a periodic QR algorithm [5] to compute all multipliers. The third method, proposed in this paper, discretizes the variational system using a multistep method and extracts the Floquet multipliers with a periodic Krylov-Schur algorithm.
To ensure a fair comparison, all methods use a total of 1024 sampling points per period. For AUTO-07p and the modified MATCONT, we set NTST = 512 and NCOL = 2.
For clarity of the figures, we plot the modulus of the computed Floquet multipliers on a logarithmic scale versus the period Fig.˜3. Around suspected bifurcation points, the Floquet multipliers change very rapidly within narrow parameter intervals. Due to the limited number of computed points, these sharp transitions are not fully resolved in our figures.
(a) First Floquet Multiplier
(b) Second Floquet Multiplier
(c) Third Floquet Multiplier
(d) Fourth Floquet Multiplier
Figure 3: Evolution of Floquet Multipliers vs. Period of the Parameterized Limit Cycle
The three largest Floquet multipliers obtained via all three methods lie on very smooth curves, indicating that these multipliers are likely quite accurate. For the largest Floquet multipliers shown in Fig.˜3, all three methods yield similar results. However, AUTO-07p fails to report the largest multiplier once it exceeds $10^{16}$. Aside from this limitation, its computed values remain consistent with other two methods. The second and third largest multipliers are nearly identical across all methods. For the smallest Floquet multiplier, both the modified MATCONT and our method indicate a monotonic decay. Notably, AUTO-07p incorrectly reports the fourth multiplier as an approximately constant value around $10^{-12}$ over a wide range of periods, while our results show a clear downward trend.
Figure 4: Error of Oscillatory Floquet Multiplier vs. Period of the Parameterized Limit Cycle
Figure˜4 shows the error in computing the oscillatory Floquet multiplier, which should be 1. Both AUTO-07p and the modified MATCONT maintain consistently high accuracy due to their use of fourth-order collocation. In comparison, our Gear4 method performs worse. The reason is as the period increases and the smallest multiplier approaches zero, the spectral gap between Floquet multipliers and parasitic eigenvalues decreases. This leads to the separation between Floquet subspace and parasitic subspaces decreases, and thus reducing accuracy of each Floquet multipliers. Although the accuracy of the oscillatory multiplier computed by our method is slightly inferior, we are still able to maintain its value close to 1. Since the primary objective is to track the evolution of all four multipliers relative to the unit circle as a parameter varies, the overall performance remains satisfactory, as it still captures all four multipliers reasonably and correctly determines stability.
6.3 Scenario 2: RF Circuits
In radio frequency (RF) circuits, accurately predicting the phase noise of oscillators is essential for evaluating the performance and stability of oscillatory systems [9]. A critical step in this prediction involves computing the perturbation projection vector (PPV) [9], which corresponds to the eigenvector associated with the dominant Floquet multiplier of an LTP system. The LTP system takes the form:
(37) $C(t)^{\mathrm{T}}\frac{\mathrm{d}y}{\mathrm{d}t}+G(t)^{\mathrm{T}}y=0,$
where $C(t)$ and $G(t)$ are $T$-periodic matrices with the following structure:
$C(t)=\begin{bmatrix}C_{11}(t)&0\\ 0&0\end{bmatrix},\quad G(t)=\begin{bmatrix}G_{11}(t)&G_{12}(t)\\ G_{21}(t)&G_{22}(t)\end{bmatrix},\quad y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}$.
Here, $C_{11}(t)$ and $G(t)$ are assumed to be nonsingular. The matrix $C_{11}(t)$ is referred to as the capacitance matrix, while $G(t)$ is the conductance matrix. The vector $y_{1}$ represents the part of the nodal voltages, and $y_{2}$ denotes the rest nodes. According to the DAE theory presented in [21], equation Eq.˜37 is classified as a semi-explicit index-1 differential-algebraic equation (DAE). It can be systematically decoupled into an ODE and an algebraic constraint:
(38) $\left\{\begin{aligned} &-\frac{\mathrm{d}}{\mathrm{d}t}y_{1}(t)+C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)\right]y_{1}=0,\\ &y_{2}(t)=-G_{22}^{-T}(t)G_{12}^{\mathrm{T}}(t)y_{1}(t).\end{aligned}\right$.
Our task is to solve the Floquet multiplier closest to 1 and its eigenvectors of the ODE part of Eq.˜38, and thus recover the PPV. The physical nature of RF circuits leads to the following three characteristics of Eq.˜38:
  1. 1.
    all Floquet multipliers have magnitudes strictly less than one, with only one or a few of them being close to $1$;
  2. 2.
    both the $C(t)$ and $G(t)$ are highly sparse;
  3. 3.
    only discrete samples of $C(t)$ and $G(t)$ are available at specific time points $0=t_{0}<t_{1}<t_{2}<\cdots<t_{p}=T$, and further sampling is not permitted.
Since MATCONT and AUTO-07p require explicit system equations to perform collocation at Gauss-Legendre points, they are not applicable in our setting where the sampling points are fixed. In contrast, our method is still applicable, and we avoid forming the explicit expression of $C_{11}^{-\mathrm{T}}(t)\left[G_{11}^{\mathrm{T}}(t)-G_{21}^{\mathrm{T}}(t)G_{22}^{-\mathrm{T}}(t)G_{12}^{\mathrm{T}}(t)\right]$.
Model Name Dimension Rank of $C(t)$ Total Sampled Points
Well Conditioned Case 18 16 448
Ill Conditioned Case 913 388 654
Table 3: PPV Cases Information
Table˜3 summarizes the basic properties of the two test cases. We use Gear4 as a benchmark and apply BE, Gear2, and Gear3 to discretize the systems, and then use the pTOAR algorithm to compute Floquet eigenpairs.
The goal is to verify as the discrete stepsize decrease, whether the convergence behavior observed in theory holds in practical, industry-scale circuits. Specifically, we examine whether the dominant eigenvector converges at the expected rate when it is well separated from the others, and whether the convergence instead occurs toward the invariant subspace when multiple Floquet multipliers cluster near 1.
(a) Floquet Multipliers Relative to the Unit Circle
(b) Separation between the Dominant Invariant Subspaces and the Remaining
(c) Dominant Eigenvector Error vs Stepsize
Figure 5: Floquet Multipliers, Subspace Separation, and Dominant Eigenvector Convergence for the Well Conditioned Case
In Fig.˜5(a), the Floquet multipliers are plotted in blue relative to the unit circle in black. Most multipliers are clustered near the origin, except for a single one lying close to 1.
The subspace separation between dominant $k$-dimension subspace (the invariant subspace of the dominant $k$ eigenvalues) and the remainder is bounded above by the spectral gap between the $k$-th and $(k+1)$-th largest eigenvalues. Consequently, the dominant one-dimensional subspace is expected to be well isolated, while this separation rapidly decay as more multipliers are included. This expectation is confirmed in Fig.˜5(b), indicating this case is a well conditioned case for solving PPV.
To verify whether the theoretical convergence rates are achieved in practice, we computed the invariant subspace for the dominant Floquet multiplier using several multistep methods with varying maximum stepsizes. As shown in Fig.˜5(c), all three methods converge at the expected theoretical rates.
In contrast, the ill-conditioned case features multiple Floquet multipliers near 1, as shown in Fig.˜6(a). Figure˜6(b) reveals that although the dominant 13-dimensional invariant subspace is well-separated from the remainder, the dominant one-dimensional subspace itself is poorly isolated. This makes accurately computing the PPV particularly challenging.
(a) Floquet Multipliers Relative to the Unit Circle
(b) Separation between the Dominant Invariant Subspaces and the Remaining
Figure 6: Floquet Multipliers, Subspace Separation for the Ill Conditioned Case
In Fig.˜7, we examine the convergence behavior of both the dominant eigenvector and the 13-dimensional subspace for the ill conditioned case. Figure˜7(a) shows that the dominant Floquet eigenvector fails to converge. In contrast, Figure˜7(b) demonstrates that the dominant 13-dimensional subspace converges at expected rate. This indicates that the clustering of the leading Floquet multipliers causes interference among their corresponding eigenvectors, degrading the convergence of any individual eigenvector. However, the collective subspace remains numerically recoverable with reasonable accuracy.
(a) Dominant Eigenvector Error vs. Maximum Stepsize
(b) Dominant 13-dim. Subspace Error vs. Maximum Stepsize
Figure 7: Convergence Behavior of Dominant Subspaces of the Ill Conditioned Case
(a) Well Conditioned Case
(b) Ill Conditioned Case
Figure 8: Computation Time vs. Step Numbers for the Well and Ill Conditioned Cases.
Figure˜8 presents the computational time required by our algorithm under different discretization methods and varying numbers of time steps. Excluding a few noticeable outliers, the results show that for a fixed discretization method, the computational time increases linearly with the number of time steps. Similarly, for a fixed number of time steps, the use of higher-order multistep methods does not lead to a significant increase in runtime. These observations are consistent with our earlier analysis of the algorithm’s computational complexity.
7 Conclusion
A multistep discretization of an LPTV system gives rise to a pPEP for computing Floquet multipliers and their invariant subspaces. This pPEP exhibits clear spectral separation: as the stepsize tends to zero, the dominant eigenvalues converge to the true Floquet multipliers, while the remaining ones decay to zero. By subspace perturbation theory, the computed multipliers and invariant subspaces converge to their exact counterparts at a rate matching the consistency order of the multistep method. Building on this, we propose the pTOAR algorithm to solve all or part of the Floquet eigenpairs efficiently. Numerical experiments demonstrate that the algorithm effectively addresses the two common scenarios, as outlined in the introduction.
Acknowledgments
We would like to acknowledge the assistance of volunteers in putting together this example manuscript and supplement.